OpenStudy (anonymous):

Find the radius of a circle with the given circumference. C = 6 sqrt 2 π in. r = _____ 3 in. 3√2 in. 6 in. 12√2 in.

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

@leon549

OpenStudy (anonymous):

r= C/2 pi

TheSmartOne (thesmartone):

\(\sf\Large C = 2\pi r\) So then trying to isolate r by itself \(\sf \Large r = \frac{C}{ 2\pi}\)

OpenStudy (anonymous):

3 sqrt 2?

TheSmartOne (thesmartone):

Correct @hhelpplzzzz

OpenStudy (anonymous):

thank you

TheSmartOne (thesmartone):

yw